# ExplainReduce: Summarising local explanations via proxies

### v1

@@ 4.1 Problem definition
(minimum loss) Given $k$, find a subset ${\bm{S}}_{c}$ with cardinality $k$ such the average loss when picking the lowest loss model from ${\bm{S}}_{c}$ is minimised, or
${\bm{S}}_{c}=\arg\min\nolimits_{{\bm{S}}\in[m]_{k}}{\left(\frac{1}{n}\sum\nolimits_{j=1}^{n}{\min\nolimits_{i\in{\bm{S}}}{\ell(g_{i}({\bf x}_{j}),{\bf y}_{j})}}\right)}.$ (5)
The final formulation is a combination of problems 1 and 2.
@@
(coverage-constrained minimum loss) Given $k$, $\varepsilon\in{\mathbb{R}}_{>0}$, and minimum coverage $c\in(0,1]$, find a $c$-covering ${\bm{S}}_{c}$ with cardinality $k$ such the average loss when picking the lowest loss model from ${\bm{S}}_{c}$ is minimised, or
${\bm{S}}_{c}=\arg\min\nolimits_{{\bm{S}}\in[m]_{k,c}}{\left(\frac{1}{n}\sum\nolimits_{j=1}^{n}{\min\nolimits_{i\in{\bm{S}}}{\ell(g_{i}({\bf x}_{j}),{\bf y}_{j})}}\right)},$ (6)
where $[m]_{k,c}=\{{\bm{S}}\in[m]_{k}\mid C({\bm{S}},\varepsilon)\geq c\}$ are the $c$-coverings of cardinality $k$.
@@ 4.2 General procedure
The ExplainReduce algorithm is outlined in Algorithm 1. Given a dataset $\mathcal{D}$, an explanation method Exp that generates a set of $m$ local explanations (a special case being one local explanation for each training data point, in which case $m=n$) and a reduction algorithm explained later in Sect. 4.3, we first use the explanation method to generate $m\leq n$ explanations for $m$ items sampled without replacement from $\mathcal{D}$. If a closed-box function is provided, we replace the original labels in $\mathcal{D}$ with predictions from $f$. We then apply the reduction algorithm to the generated set of local explanations $[m]$ and receive the proxy set ${\bm{S}}_{c}\subseteq[m]$. Finally, for each sampled data item, we pick the local explanation in the proxy set ${\bm{S}}_{c}$ with minimal loss, mapping the items and the proxies.
@@
Input: $\mathcal{D}\leftarrow\{(\bm{x}_{i},\hat{\bm{y}}_{i})|i\in[n]\}$: dataset; ${\bm{G}}$: the set of $m$ local explanations; reduce: method to find ${\bm{S}}_{c}$, parametrised optionally by $\varepsilon$, $c$, or $k$, see Sect. 4.3.
@@
     $\bm{S}_{c}\leftarrow\textrm{reduce}(\bm{G},\varepsilon,c,k)$triangleright reduce is defined in Sect. 4.3
     $\text{map}\leftarrow\{\}$triangleright a mapping between items in $\mathcal{D}$ and the local models
@@
4.3 Reduction algorithms
In this section, we briefly overview the practical implementations of the reduction algorithms (function reduce in Alg. 1) used to solve the problems outlined in the previous section.
max coverage: Problem 1 is a variant of the NP-complete partial set covering problem, sometimes called max $k$-cover. The equivalence is obvious if we consider an item $(\bm{x}_{j},\bm{y}_{j})\in\mathcal{D}$ covered by model $g_{i}$ if $\ell(g_{i}(\bm{x}_{j}),\bm{y}_{j})\leq\varepsilon$. We solve Prob. 1 exactly using integer programming (implemented by the pulp Python library [23]) and approximately using a greedy algorithm. In the greedy approach, given fixed $k$, we iteratively pick local models $g\in\bm{G}$ such that the marginal increase in coverage is maximised with each iteration until $k$ models have been chosen. Our problem is submodular, as adding each new model to the subset cannot decrease the coverage. It has been shown that in this case, the greedy algorithm has a guaranteed lower bound to achieve coverage at least $1-\left((k-1)/k\right)^{k}$ times the optimal solution [24].
Notably, this approximation ratio only applies to the original set of data items and their associated local explanations. If we apply the proxy sets to novel data, we should use standard machine learning tools — such as a separate validation set — to ensure that the model performs appropriately.
min loss: Problem 2 is an example of a supermodular minimisation problem. Let $f(\bm{S})=\left(\frac{1}{n}\sum_{j=1}^{n}{\min_{i\in{\bm{S}}}{\ell(g_{i}({\bf x}_{j}),{\bf y}_{j})}}\right)$ and $\bm{A},\bm{B}:\;\bm{A}\subset\bm{B}\subset\bm{G}$ be subsets of local models. Additionally, let $v$ be a local model not contained in $\bm{B}$. Clearly, the decrease in loss by adding $v$ to the larger set $\bm{B}$ must be, at most, as great as adding the same model $v$ to $\bm{A}$. In other words,
$f(\bm{A}\cup\{v\})-f(\bm{A})\leq f(\bm{B}\cup\{v\})-f(\bm{B})\quad\forall\bm{A}\subset\bm{B}\subset\bm{G},\;v\notin\bm{B},$ (7)
which is the definition of supermodularity [25].
Supermodular minimisation problems are known to be NP-hard. Hence, we only use a greedy ascent algorithm to solve Problem 2, as finding an exact solution is computationally expensive due to the continuous nature of loss. In the worst case, the search for the optimal subset would require $\binom{m}{k}$ comparisons. Like above, we iteratively pick local models with the best possible decrease in marginal loss until we reach $k$ chosen local models. In the general case, a multiplicative approximation ratio for a supermodular minimisation problem may not exist due to the optimal solution having a value $f(\bm{S}^{*})=0$, while the greedy algorithm may converge to a solution with non-zero loss. However, in our case, finding a subset of models with exact zero loss for all items is unlikely. In [26], the author derives a curvature-based approximation ratio for a greedy descent (worst out) algorithm. Based on this analysis, authors of [27] derive an approximation ratio for greedy ascent for probabilistic decision trees. Neither of these approaches is directly applicable to our setting, and hence, we cannot give a closed-form approximation ratio. However, as we show in Section 5.4, the empirical approximation ratios for the greedy ascent algorithm are reasonable. Moreover, we show how the greedy approximation performs nearly equally, if not better, on unseen data compared to the exact solution for our datasets.
const min loss: We solve Problem 3 again with a greedy ascent algorithm, i.e., iteratively picking the best model to include in the subset ${\bm{S}}_{c}$ until the maximum coverage has been met. The only difference to min loss algorithm is how the models are scored. If the coverage constraint has been met, we simply follow the procedure in min loss and select the model with the best possible marginal loss decrease. Otherwise, we divide the marginal loss decrease $\Delta\ell_{i}$ we would get by including model $i$ in the proxy set with the marginal coverage increase of including the same model $\Delta c_{i}$. This reduction method can be implemented with a hard constraint by having the algorithm throw an exception if the coverage is not met or with a soft constraint where the resulting proxy set is returned regardless of satisfying the constraint. In this paper, we opt to use the softly constrained variant exclusively. We can see that in the case where the coverage constraint is not met, Problem 3 with this scoring scheme is also a supermodular minimisation problem and equivalent to Problem 2 with a different optimisation objective, namely the coverage-normalised loss. Similar to Problem 2, we cannot give an exact approximation ratio for this algorithm, but the empirical results in Section 5.4 suggest good performance.
clustering: Above, we have treated the problem as a set covering problem. We can also approach the problem from the unsupervised clustering perspective: can we cluster the training data to find a proxy set? Given the number of clusters $k$, we can cluster the training data items $\bm{X}$, the local model parameters $\bm{B}$, or the training loss matrix $\bm{L}$ using some clustering algorithm. In this paper, we use k-means with Euclidean distance for $\bm{X}$ and $\bm{L}$, and cosine distance for the local model parameters $\bm{B}$. After performing the clustering, we pick the local model closest to the cluster centroid for each cluster to form the proxy set.
@@ 4.4 Performance measures
Fidelity Fidelity [4] measures the adherence of a surrogate model to the complex closed-box function. Given a closed-box function $f$, we fidelity is the loss between the closed-box model prediction $\hat{\bm{y}}=f(\bm{x})$ and the surrogate model prediction:
$\textrm{fidelity}=\frac{1}{n}\sum\nolimits_{i=1}^{n}\ell(g_{i}(\bm{x}_{i}),\hat{y}_{i}),$ (8)
where $g_{i}$ is the relevant local model either from the full set of local explanations or the proxy model set.
@@
Instability Instability [4] (sometimes also referred to as stability, despite lower values denoting better performance) measures how much a slight change in the input changes the explanation. In practice, we model the slight change by measuring the loss of a given local model $g_{i}$ associated with item $\bm{x}_{i}$ with its $\kappa$ nearest neighbours:
$\textrm{instability}=\frac{1}{n}\sum\nolimits_{i=1}^{n}\frac{1}{\kappa}\sum\nolimits_{j\in\textrm{NN}_{\kappa}(i)}\ell(g_{i}(\bm{x}_{j}),y_{j}).$ (9)
This paper uses a fixed number of $\kappa=5$ nearest neighbours when we report instability values ($\left|\textrm{NN}_{5}(i)\right|=5$).
5 Experiments
In this section, we demonstrate the performance of the proxy sets produced via different XAI methods and reduction strategies. We experiment on a variety of different datasets, consisting of both classification and regression tasks. For each of the datasets used, we train a closed-box model and produce the initial set of local explanations using that model. A brief description of the datasets can be found in Table 1, and further details in Appendix A.
@@ 5 Experiments
Synthetic 5000 × 11 Regression Random Forest [5]
Air Quality 7355 × 12 Regression Random Forest [28]
Life Expectancy 2938 × 22 Regression Neural Network [29]
Vehicle 2059 × 12 Regression SVR [30]
Gas Turbine 36733 × 9 Regression Adaboost [31]
QM9 133766 × 27 Regression Neural Network [32]
Jets 266421 × 7 Classification Random Forest [33]
HIGGS 100000 × 28 Classification Gradient Boosting [34]
Table 1: Summary of datasets used in experiments.
In the experiments, we measure fidelity mostly with test data, i.e., data not used in generating the full set of local explanations. This metric better captures the adherence of the explainer to the closed-box model than calculating the same value on the data used to generate the local explanations.
The local explanation methods discussed in this paper are generally not generative, that is, there is no obvious mapping from a new covariate ${\bm{x}}$ to an explanation model in ${\bm{S}}_{c}$. Therefore, to use either the full explanation sets or the proxies to produce predictions, we need to generate another mapping between novel items and local models in the explanation set. The simplest of such mappings is to use the distance in the data space: $\Phi_{\textrm{new}}(\bm{x})=\arg\min\nolimits_{i\in{\bm{S}}_{c}}||\bm{x}_{i}-\bm{x}||^{p},\;i\in(1,...,n)$, where $||a||^{p}$ is the $p$-norm. This new mapping is used for unlabelled items only and is in addition to the loss-minimising mapping used in the reduction process to find an initial mapping with the “training” data (data used to generate the local explanations) described in Section 4.2.
@@ 5.1 Case studies
We begin our examination by demonstrating the ExplainReduce procedure with two case studies to give the reader a better intuition of the procedure and its possible usage.
@@ 5.1.1 Synthetic data
To give a simple example of the procedure, we apply it to a synthetic dataset. The dataset, which is described in detail in Appendix A, is generated by producing $k=4$ clusters in the input space and generating a random, different linear model for each cluster. We then generate labels by applying a local model to the data items based on their cluster ID and adding Gaussian noise. The dataset is then randomly split into a training set and a test set, and we train a smoothgrad explainer on the training data. For reduction, we use a greedy max coverage algorithm, where $\varepsilon$ is defined as the 10th percentile of the loss matrix $\bm{L}$.
In Fig. 2, we show a PCA of the items in the test set on the left, coloured with the ground truth cluster labels, and how the test items are mapped to proxy models on the right. Overall, we can see that most items get mapped to the correct proxy model for that particular cluster. The small impurity in the clusters stems both from the Gaussian noise and the approximative nature of the greedy coverage-maximising algorithm. Furthermore, as Fig. 3 shows, the reduced proxy models (red) correspond well with the ground truth models (blue). The proxy set thus serves well as a generative global explanation for the dataset.
Figure 2: Ground PCA of a synthetic test dataset (left) and the same dataset where colours correspond to reduced model indices, based on smoothgrad local models and reduced using a greedy coverage-maximising algorithm (right). We can see that the reduction is able to faithfully approximate the ground truth clustering.
Figure 3: Radar plots showing ground truth local model parameters (blue) and corresponding proxy model parameters (red). The reduction method is able to find very similar surrogate models to the ground truth.
@@ 5.1.2 Particle jet classification
In a previous work [3], we analysed a dataset containing simulated LHC proton-proton collisions using slisemap. These collisions can create either quarks or gluons, which decay into cascades of stable particles called jets. These jets can then be detected, and we can train the classifier to distinguish between jets created by quarks and gluons based on the jet’s properties.
@@
We first trained a random forest model on the data and then applied slisemap to find local explanations. We clustered the slisemap local explanations to 5 clusters and analysed the average models for each cluster, and found them to adhere well to physical theory. When we generate 500 slisemap local explanations on the same dataset and apply the const min loss reduction algorithm with $k=4$ proxies, we find the proxy set depicted in Fig. 4. The left panel shows a swarm plot where each item is depicted on a horizontal line based on the random forest-predicted probability of the jet corresponding to a gluon jet and coloured based on which proxy model is associated with the item. The right panel shows the coefficients of the proxy models, which are regularised logistic regression models. We find that the proxies are similar to the cluster mean models shown in the previous publication and show similar adherence to the underlying quantum chromodynamic theory [35]. For example, wider jets (high jetGirth and QG_axis2) are generally more gluon-like, and therefore these parameters are essential in classifying the jets. Incidentally, proxy 3 (red) is associated with the most quark- and gluon-like jets and has high positive coefficients for both parameters. Similarly, proxy 0 (blue) has negative coefficients for momentum (jetPt) and QG_ptD, which measures the degree to which the total momentum parallels the jet. Both of these features indicate a more quark-like jet.
@@
Figure 4: Example of the ExplainReduce procedure on a particle jet classification task. The dataset consists of LHC proton-proton collision particle jets that are created by decaying gluons or quarks. We train a random forest classifier on the dataset and use slisemap to generate 500 explanations, which are then reduced to $k=4$ proxies using the const min loss reduction algorithm. The left panel shows a swarm plot of the 500 items sorted horizontally based on the predicted probability of corresponding to a gluon; the y-axis has no significance. The right panel shows the coefficients of the logistic regression proxy models, which match the underlying physical theory.
5.2 Fidelity of proxy sets

### v2

@@ 4.1 Problem definition
(Min Loss) Given $k$, find a subset ${\bm{S}}_{c}$ with cardinality $k$ such that the average loss $\mathcal{L}(\bm{S})$ when picking the lowest loss model from ${\bm{S}}_{c}$ is minimised, or
${\bm{S}}_{c}$ $=\arg\min\nolimits_{{\bm{S}}\in[m]_{k}}\mathcal{L}(\bm{S})\quad\text{where}$ (4)
$\mathcal{L}(\bm{S})$ $=\left(\sum\nolimits_{j=1}^{n}{\min\nolimits_{i\in{\bm{S}}}{\ell(g_{i}(\bm{x}_{j}),\bm{\hat{y}}_{j})}}/n\right).$ (5)
While minimizing $\mathcal{L}(\bm{S})$ is a supermodular minimization problem, it can be framed as the maximization of a non-negative, monotone submodular function $f(\bm{S})=\mathcal{L}_{base}-\mathcal{L}(\bm{S})$, where $\mathcal{L}_{base}$ represents the average loss of an initial reference point (e.g., the loss of an empty set or a single starting model). In this framework, $f(\bm{S})$ quantifies the total reduction in loss achieved by the proxy set.
By applying the greedy algorithm (reduce) to this improvement function, we leverage the classic result by [nemhauser1978analysis]: the greedy solution $\bm{S}_{g}$ is guaranteed to capture at least $(1-1/e)\approx 63.2\%$ of the total improvement possible by the optimal subset $\bm{S}^{*}$ of the same size. Formally, $\mathcal{L}_{base}-\mathcal{L}(\bm{S}_{g})\geq(1-1/e)(\mathcal{L}_{base}-\mathcal{L}(\bm{S}^{*}))$. In practice, we can solve the problem using as marginal gain metric the marginal increase in negative loss:
$\Delta\mathcal{L}(i|\bm{S})=\mathcal{L}(\bm{S})-\mathcal{L}(\bm{S}\cup\{i\}).$ (6)
We refer to this algorithm with Min Loss.
4.1.3 Joint utility maximisation.
While Problems 1 and 2 address coverage and faithfulness individually, a practical proxy set must balance both. We propose a third formulation that treats the proxy selection task as multi-objective optimisation problem.
@@
(Balanced) Given $k$ and $\varepsilon\in{\mathbb{R}}_{>0}$ and a tradeoff parameter $\lambda\in[0,1]$, find a subset $\bm{S}_{c}$ of cardinality $k$ such the joint utility of coverage and normalised loss reduction, $U(\bm{S})$, is maximised, or
${\bm{S}}_{c}$ $=\arg\max\nolimits_{{\bm{S}}\in[m]_{k}}U(\bm{S})\quad\text{where}$ (7)
$U(\bm{S})$ $=\lambda\cdot C(\bm{S},\varepsilon)+(1-\lambda)\cdot\left(\frac{\mathcal{L}_{base}-\mathcal{L}(\bm{S})}{\mathcal{L}_{base}}\right),$ (8)
where $\mathcal{L}_{base}$ is the initial average loss (e.g., the loss of the data-generating process or a single-model baseline).
Since $U(\bm{S})$ is a linear combination of two monotone submodular functions, $U(\bm{S})$ is itself monotone submodular. Consequently, solving Problem 3 via Alg. 1 provides constant-factor approximation guarantee. Specifically, the greedy solution $\bm{S}_{g}$ achieves utility such that $U(\bm{S}_{g})\geq(1-1/e)U(\bm{S}^{*})$, where $\bm{S}^{*}$ is the optimal subset of size $k$. In this paper, we use $\lambda=0.5$ to balance coverage and fidelity; in our tests (Appendix E), the results were not overly sensitive to $\lambda$, making $\lambda=0.5$ a straightforward choice. We refer to this algorithm with Balanced.
While we have presented worst-case guarantees for our approaches to solving Problems 1-3, in our experiments, the performance of these greedy algorithms is close to the true optimum (Appendix G).
@@ 4.2 General procedure
The ExplainReduce algorithm is outlined in Algorithm 2. The algorithm works in three stages. First, a large set of local explanations $\bm{G}$ is generated using an XAI method on the dataset $\mathcal{D}$ and a closed-box model $f$. Second, we use an optimisation algorithm reduce to find an approximately optimal proxy set $\bm{S}$. The options for the optimisation algorithm are described in Section 4.1 above. Finally, each data item in $\mathcal{D}$ is mapped to a proxy model in $S$ that best approximates the closed-box model for that item, i.e., has the lowest loss. The algorithm outputs the optimal proxy set $\bm{S}_{c}$ and the aforementioned mapping.
@@
Input: $\mathcal{D}\leftarrow\{(\bm{x}_{i},\hat{\bm{y}}_{i})|i\in[n]\}$: dataset; $\bm{G}$}: the set of $m$ local explanations; reduce: method to find ${\bm{S}}_{c}$, parametrised optionally by $\varepsilon$, $c$, or $k$, see Sect. 4.1.
@@
  $\bm{S}_{c}\leftarrow\textrm{reduce}(\bm{G},\varepsilon,c,k)$ $\triangleright$ reduce is defined in Sect. 4.1
  $\text{map}\leftarrow\{\}$ $\triangleright$ a mapping between items in $\mathcal{D}$ and the local models
@@ 4.3 Performance measures
Fidelity Fidelity [guidotti2018survey] measures the adherence of a surrogate model to the complex closed-box function. Given a closed-box function $f$, we fidelity is the loss between the closed-box model prediction $\hat{\bm{y}}=f(\bm{x})$ and the surrogate model prediction:
$\textrm{fidelity}=\sum\nolimits_{i=1}^{n}\ell(g_{i}(\bm{x}_{i}),\hat{y}_{i})/n,$ (9)
where $g_{i}$ is the relevant local model, either from the full set of local explanations or the proxy model set. Somewhat paradoxically, lower fidelity is better. In this paper, we use mean square error as the loss function for regression tasks. For classification tasks, we use the squared Hellinger distance of class probabilities,
$\ell_{c}(\hat{\bm{y}},\tilde{\bm{y}})=\sum\nolimits_{i=1}^{n}\left\|\sqrt{\hat{\bm{y}}_{i}}-\sqrt{\tilde{\bm{y}}_{i}}\right\|^{2}_{2}/2,$ (10)
where $\tilde{\bm{y}_{i}}=g_{i}(\bm{x}_{i})\in\Delta^{p-1}$, where $\Delta^{p-1}$ is the p-simplex containing class probabilities of the $p$ classes; a more numerically stable replacement for the Kullback-Leibler divergence as in bjorklund2023slisemap.
@@
Instability Instability [guidotti2018survey] (sometimes also referred to as stability, despite lower values denoting better performance) measures how much a slight change in the input changes the explanation. In practice, we model the slight change by measuring the loss of a given local model $g_{i}$ associated with item $\bm{x}_{i}$ with its $\kappa$ nearest neighbours:
$\textrm{instability}=\sum\nolimits_{i=1}^{n}\sum\nolimits_{j\in\textrm{NN}_{\kappa}(i)}\ell(g_{i}(\bm{x}_{j}),\hat{y}_{j})/(n\,\kappa).$ (11)
In this work, we use a fixed number of $\kappa=5$ nearest neighbours to report instability values ($\left|\textrm{NN}_{5}(i)\right|=5$).
5 Results
In this section, we evaluate ExplainReduce on a diverse set of regression and classification tasks and XAI methods to assess its ability to produce compact, interpretable global explanations without sacrificing fidelity. Our experiments seek to address three key questions: (i) how well do proxy sets approximate the original closed-box model compared to the full set of local explanations, (ii) how does performance vary with the proxy set size and the number of initial local explanations, and (iii) how does ExplainReduce compare to existing aggregation methods.
We run experiments on twelve datasets containing both synthetic and real-world examples (Table 1, Appendix B). For each dataset, we train a closed-box model, generate local explanations using multiple XAI methods and apply ExplainReduce to find proxy sets. The closed-box models have been chosen based on task-specific performance. In cases where multiple model types showed similar results, closed-box models were chosen to maximise model type diversity. Performance is measured primarily on test data—items not used to generate the initial set of local explanations—to best capture general adherence to the closed-box model. In this manner we can measure whether the explanation set contains a reasonable explanation for any item.
@@ 5 Results
Synthetic 5000 × 11 Regression Random Forest [bjorklund2023slisemap]
Air Quality 7355 × 12 Regression Random Forest [oikarinenDetectingVirtualConcept2021]
Life Expectancy 2938 × 22 Regression Neural Network [rajarshi2017life]
Vehicle 2059 × 12 Regression SVR [birla2022vehicles]
Gas Turbine 36733 × 9 Regression Adaboost [2019gas]
QM9 133766 × 27 Regression Neural Network [ramakrishnan2014Quantum]
Adult 48842 × 14 Classification Neural Network [becker1996adult]
Churn 5000 × 20 Classification Random Forest [blake1998churn]
HIGGS 100000 × 28 Classification Gradient Boosting [whiteson2014higgs]
Jets 266421 × 7 Classification Random Forest [CMS:opendata]
Spam 4601 × 57 Classification Gradient Boosting [hopkins1999spam]
Telescope 13376 × 10 Classification Neural Network [bock2004telescope]
Table 1: Summary of datasets used in experiments. Detailed dataset descriptions available in Appendix B.
A practical challenge is that local explanation sets do not have an obvious mapping from unseen items to explanations. To address this, we adopt a simple nearest-neighbour approach: each novel item is assigned to a proxy corresponding to the training item that is closest in the feature space using Euclidean distance. While using this approach has obvious limitations, we nonetheless found it to work reasonably well, as will be demonstrated. In practical implementations, when using ExplainReduce with novel items, we recommend that users use whatever distance measure is most fitting for their data.
The following subsections present case studies to build intuition, followed by systematic comparisons of reduction algorithms and baselines.
@@ 5.1.1 Synthetic data
To give a simple example of the procedure, we apply it to a synthetic dataset. The dataset, which is described in detail in Appendix B, is generated by producing $k=4$ clusters in the input space and generating a random, different linear model for each cluster. We then generate labels by applying a local model to the data items based on their cluster ID and adding Gaussian noise. The dataset is then randomly split into a training set and a test set, and we train a smoothgrad explainer on the training data. For reduction, we use a greedy max coverage algorithm, where $\varepsilon$ is defined as the 10th percentile of the loss matrix $\bm{L}$.
In Fig. 2, we show a PCA of the items in the test set on the left, coloured with the ground truth cluster labels, and how the test items are mapped to proxy models on the right. Overall, we can see that most items get mapped to the correct proxy model for that particular cluster. The small impurity in the clusters stems from both the Gaussian noise and the approximative nature of the greedy coverage-maximising algorithm. Furthermore, as Fig. 3 shows, the reduced proxy models (red) correspond well with the ground truth models (blue). The proxy set thus serves well as a generative global explanation for the dataset.
Figure 2: ExplainReduce is able to correctly associate items with the correct local explanation. The left panel shows a PCA embedding of a synthetic test dataset with ground truth cluster labels, and the right panel shows the same dataset where colours correspond to reduced model indices. The XAI method used here is smoothgrad, and the max coverage algorithm was used for the reduction. We can see that the reduction faithfully approximates the ground truth clustering.
Figure 3: The proxy models found by ExplainReduce (red) correspond well to ground truth local models (blue) used to generate a synthetic dataset. The spokes in the radar plots correspond to linear model coefficients associated with each feature.
@@ 5.1.2 Particle jet classification
In a previous work [seppalainen2023using], we analysed a dataset containing simulated LHC proton-proton collisions using slisemap. These collisions can create either quarks or gluons, which decay into cascades of stable particles called jets. These jets can then be detected, and we can train the classifier to distinguish between jets created by quarks and gluons based on the jets’ properties.
@@
We first trained a random forest model on the data and then applied slisemap to find local explanations. We clustered the slisemap local explanations into 5 clusters, analysed the average models for each cluster, and found them to adhere well to physical theory. When we generate 500 slisemap local explanations on the same dataset and apply ExplainReduce with the Balanced reduction algorithm with $k=4$ proxies, we find the proxy set depicted in Fig. 4. The left panel shows a swarm plot where each item is depicted on a horizontal line based on the probability predicted by the random forest for a jet corresponding to a gluon jet. It is coloured based on which proxy model is associated with the item. The right panel shows the coefficients of the regularised logistic regression proxy models. We find that the proxies are similar to the cluster mean models shown in the previous publication and show similar adherence to the underlying quantum chromodynamic theory [cms2013performance]. For example, wider jets (high jetGirth and QG_axis2) are generally more gluon-like, and therefore these parameters are essential in classifying the jets. Incidentally, proxy 3 (red) is associated with the most quark- and gluon-like jets and has high positive coefficients for both parameters. Similarly, proxy 0 (blue) has negative coefficients for momentum (jetPt) and QG_ptD, which measures the degree to which the total momentum parallels the jet. Both of these features indicate a more quark-like jet.
@@
Figure 4: ExplainReduce is able to find proxy sets that adhere to the underlying physical theory. The dataset consists of LHC proton-proton collision particle jets that are created by decaying gluons or quarks. We train a random forest classifier on the dataset and use slisemap to generate 500 explanations, which are then reduced to $k=4$ proxies using the Balanced reduction algorithm. The left panel shows a swarm plot of the 500 items sorted horizontally based on the predicted probability of corresponding to a gluon; the y-axis has no significance. The right panel shows the coefficients of the logistic regression proxy models, which match particle-physics-based intuition.
5.2 Comparison of reduction algorithms
5.2.1 Fidelity of proxy sets